# Tyche: Stochastic In-Context Learning for Medical Image Segmentation

### v1

@@ 1 Introduction
Segmentation is a core step in medical image analysis, for both research and clinical applications. However, current approaches to medical image segmentation fall short in two key areas. First, segmentation typically involves training a new model for each new modality and biomedical domain, which quickly becomes infeasible given the resources and expertise available in biomedical research and clinical environments. Second, models most often provide a single solution, whereas in many cases, the target image contains ambiguous regions, and there isn’t a single correct segmentation. This ambiguity can arise from noisy or low contrast images, variation in the task definition, or human raters’ interpretations and downstream goals [12, 55]. Failure to take this ambiguity into account can affect downstream analysis, diagnosis, and treatment.
Recent work tackles these issues separately. In-context learning methods generalize to unseen medical image segmentation tasks, employing an input context or prompt to guide inference [19, 129, 128]. These methods are deterministic and predict a single segmentation for a given input image and task.
Separately, stochastic or probabilistic segmentation methods output multiple plausible segmentations at inference, reflecting the task uncertainty [11, 67, 95]. Each such model is trained for a specific task, and can only output multiple plausible segmentations at inference for that task. Training or fine-tuning a model for a new task requires technical expertise and computational resources that are often unavailable in biomedical settings.
@@ 2 Related Work
Biomedical segmentation is a widely-studied problem, with recent methods dominated by UNet-like architectures [107, 6, 51]. These models tackle a wide variety of tasks, such as different anatomical regions, different structures to segment within a region, different image modalities, and different image settings. With most methods, a new model has to be trained or fine-tuned for each combination of these. Additionally, most models don’t take into account image ambiguity, and provide a single deterministic output.
Multiple Predictions. Uncertainty estimation can help users decide how much faith to put in a segmentation [26] and guide downstream tasks. Uncertainty is often categorized into aleatoric, uncertainty in the data, and epistemic, uncertainty in the model [29, 61]. In this work, we focus on aleatoric uncertainty. Medical images are also heteroscedastic in that the degree of uncertainty varies across the image.
Different strategies exist to capture uncertainty. One can assign a probability to each pixel [62, 45, 52, 76], or use contour strategies and difference loss functions to predict the largest and smallest plausible segmentations [73, 132]. These strategies however do not capture the correlations across pixels. To address this, some methods generate multiple plausible label maps given an image [95, 11, 67, 68, 130]. To achieve this, one can directly model pixel correlations, such as through a multivariate Gaussian distribution (with low rank) covariance [95], or more complex distributions [15]. Alternatively, various frameworks combine potentially hierarchical representations for UNet-like architectures with variational auto-encoders [67, 11, 68]. More recently, diffusion models have been used for ensembling [130] or to produce stochastic segmentations [105, 133]. Some methods explicitly model the different annotators to capture ambiguity [48, 100, 111, 124]. But these methods do not apply to our framework where the number of annotators and their characteristics are unknown.
@@
In-context Learning. Few-Shot frameworks use a small set of examples to generalize to new tasks [31, 80, 98, 117, 101, 113, 134], sometimes by fine-tuning an existing pretrained model [32, 99, 125, 120]. In-context learning segmentation methods (ICL) use a small set of examples directly as input to infer label maps for a task [19, 9, 63, 129, 129, 63]. This enables them to generalize to new tasks. For example, UniverSeg uses an enhanced UNet-based architecture to generalize to medical image segmentation tasks unseen during training [19]. We build on these ideas to enable segmentation of new tasks without the need to re-train, but expand this paradigm to model stochastic segmentations.
Test Time Augmentation. The test-time augmentation (TTA) strategy uses perturbations of a test input and ensembles the resulting predictions. Existing TTA frameworks model accuracy [116, 33, 119, 64], robustness [25], and estimates of uncertainty [90, 5]. Test-time augmentation has been applied to diverse anatomies and modalities including brain MRI and retinal fundus [3, 5, 49, 97, 53, 127]. Prior work has formalized the variance of a model’s predictions over a set of input transformations as capturing aleatoric uncertainty [126, 127, 5].
@@ 3.1.1 Neural Network
We use a convolutional architecture focused on interacting representations of sets of flexible sizes using a modified version of the usual UNet structure [107].
@@
SetBlock. We introduce a new operation called SetBlock, which interacts the candidate representations $U=\{u_{i}\}_{i=1}^{K}$, with the context representations $V=\{v_{i}\}_{i=1}^{S}$, illustrated in Figure 4. We use the CrossBlock [19] as a building block for this new layer. The $\text{CrossBlock}(u,V)\rightarrow(u^{\prime},V^{\prime})$ compares an image representation $u$ to a context set representation $V$ through convolutional and averaging operations, and outputs a new image representation $u^{\prime}$ and a new set representation $V^{\prime}$ (Figure 3). $\text{SetBlock}(U,V)\rightarrow(U^{\prime},V^{\prime})$ builds on CrossBlock and performs a set to set interaction of the entries of $U$ and $V$:
@@
[…]
@@ 3.1.2 Best candidate Loss
where $y_{r}^{t}$ is a segmentation from rater $r$, and $\mathcal{L}_{Dice}$ is a weighted sum of soft Dice loss [94] and categorical crossentropy. By only back-propagating through the best prediction among $K$ candidates, the network is encouraged to produce diverse solutions [22, 41, 66, 81].
@@ 3.1.3 Training Data
We employ a large dataset of single- and multi-rater segmentations across diverse biomedical domains. We then use data augmentation [19], as described in B.3.
@@ 4.1 Data
Single-Annotator Data. For single annotator data, we build on MegaMedical used in recent publications [19, 131] and employ a collection of 73 datasets, of different public biomedical domains and different modalities [14, 1, 56, 77, 39, 59, 79, 86, 103, 122, 110, 115, 104, 138, 84, 139, 83, 19, 35, 137, 21, 34, 24, 16, 17, 42, 54, 57, 74, 75, 82, 78, 85, 118, 87, 121, 106, 69, 96, 7, 36, 43, 71, 72, 88, 89, 91, 112, 40, 50, 2, 135]. MegaMedical spans a variety of anatomies and modalities, including brain MRI, cardiac ultrasound, thoracic CT and dental X-ray. We also use synthetic data involving simulated shapes, intensities, and image artifacts [19, 44]. The single-annotator datasets used for out-of-domain (O.D.) testing are: PanDental [1], WBC [139], SCD [104], ACDC [14], and SpineWeb [138].
Multi-Annotator Data. For multi-annotator I.D. data, we use four datasets from Qubiq [92]: Brain Growth, Brain Lesions, Pancreas Lesions, and Kidney. We also simulate a multi-rater dataset consisting of random shapes (blobs). For the O.D. multi-annotator data we use four datasets. One contains hippocampus segmentation maps on brain MRIs from a large hospital. We crop the volumes around the hippocampus [67] to focus on the areas where the raters disagree. The second is a publicly available lung nodule dataset, LIDC-IDRI [4]. This dataset is notable for the substantial inter-rater variability. It contains 1018 thoracic CT scans, each annotated by 4 annotators from a pool of 12 annotators. Finally, we also use retinal fundus images, STARE [47], annotated by 2 raters, and prostate data from the MICCAI 2021 QUBIQ challenge [92], annotated by 6 raters on two tasks. Single and multi-annotator combined, our O.D. group contains 20 tasks unseen at training time (some datasets have several tasks).
@@ 4.2 Evaluation
In the multi-annotator setting, we evaluate using both best candidate Dice score, also called maximum Dice score, as well as Generalized Energy Distance (GED)[13, 109, 123]. GED is commonly used in the stochastic segmentation literature to asses the difference between the distribution of predictions and the distribution of annotations [67, 11, 95, 105, 133]. GED has limitations, such as rewarding excessive prediction diversity [105]. Let $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ be the set of annotations, GED is defined as:
@@
where $p,p^{\prime}\sim\mathcal{Y}$, $\hat{p},\hat{p}^{\prime}\sim\hat{\mathcal{Y}}$ and $d(\cdot,\cdot)$ is a distance metric. We use the Dice score [30].
@@ 4.3 Benchmarks
In-Context Methods. We compare to deterministic frameworks that can leverage a context set: a few-shot method, SENet [108], and two in-context learning (ICL) methods, UniverSeg [19] and SegGPT [129]. We train UniverSeg and SENet with the same data split strategies and the same sets of augmentation transforms as for Tyche. For SegGPT, we use the public model, trained on a mix of natural and medical images. Figure 13 in the Supplemental Material shows that UniverSeg trained with additional data outperforms its public version.
Stochastic Upper Bounds. We compare to task-specialized probabilistic segmentation methods that are trained-on and perform well on specific datasets. We independently train Probabilistic UNet [67], PhiSeg[11] and CIDM, a recent diffusion network [105], on each of the 20 held-out tasks. For each task, we train three model variants: no augmentation, weak augmentation, and as much augmentation as for the Tyche targets. For each benchmark variant, we train on a O.D. development split and select the model that performs best on the corresponding O.D. validation split. We then compare these benchmarks to Tyche on the held-out O.D. test splits.
@@
Interactive Segmentation Methods. We compare to two interactive methods: SAM [66] and SAM-Med2D [23]. These methods can provide multiple segmentations, but, unlike Tyche, require human interaction, which is outside our scope. SAM also has a functionality to segment all elements in an image, however less optimized for medical imaging. We assume that the SAM-based models have access to the same information as the ICL methods: several image-segmentation pairs as context to guide the segmentation task. We fine-tune SAM using our I.D. development datasets. To replace the human interaction, we provide a bounding box, the average context label map, and 10 clicks, 5 positive and 5 negative as input. With SAM-Med2D, we use a bounding box, and several positive and negative clicks as input. For both SAM and SAM-Med2D, we generate clicks and bounding box from the average context label map.
@@
SegGPT $0.10\pm 0.01$ $0.68\pm 0.01$ $0.94\pm 0.01$ $0.89\pm 0.03$ $0.02\pm 0.01$
In-Context SENet $0.68\pm 0.01$ $00.00\pm 0$ 0.00 $0.83\pm 0.02$ $0.83\pm 0.02$ $0.30\pm 0.03$
@@ 4.4 Experiments
Inference Setting. We use a fixed context of 16 image-segmentation pairs, because existing in-context learning systems show minimal improvements beyond this size [19]. Because there is variability in performance depending on the context sampled, we sample 5 different context sets for each datapoint and average performance. Similarly, for the stochastic upper bounds and interactive methods, we do 5 rounds of sampling $K_{i}$ samples.
@@ 5.3 Inference Runtime
UniverSeg $96.62\pm 0.61$ $1.2$M
SegGPT $2857.19\pm 4.38$ $370$M
SENet $14.91\pm 0.21$ $0.89$M
FT-SAM $1036.75\pm 4.61$ $94$M
SAM-Med2D $188.8\pm 7.58$ $91$M
PhiSeg $11.35\pm 0.672$ $21.1$M
ProbaUNet $8.44\pm 0.46$ $5$M
CIDM $1.7\times 10^{5}\pm 2748$ $85.6$M
Tyche-IS $128.57\pm 2.626$ $1.2$M
Tyche-TS $18.09\pm 0.61$ $1.7$M
@@ B.1 Medical Image Data
Megamedical. We build on the dataset collection proposed in [19, 131], using the similar preprocessing methods. The complete list of datasets is presented Table 10.
@@
Synthetic Data. We use a set of synthetic tasks to enhance the performance of our networks, similar to [19]. Some examples are shown in Figure 11.
@@ B.2 Training of Tyche-TS
translate$\in[0,0.1]$
Random Affine 0.25 scale$\in[0.9,1.1]$
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast$\in[0.5,1.5]$
@@
translate$\in[0,0.2]$
Random Affine 0.5 scale$\in[0.8,1.1]$
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast$\in[0.8,1.2]$

### v2

@@ 1 Introduction
Segmentation is a core step in medical image analysis, for both research and clinical applications. However, current approaches to medical image segmentation fall short in two key areas. First, segmentation typically involves training a new model for each new modality and biomedical domain, which quickly becomes infeasible given the resources and expertise available in biomedical research and clinical environments. Second, models most often provide a single solution, whereas in many cases, the target image contains ambiguous regions, and there isn’t a single correct segmentation. This ambiguity can arise from noisy or low contrast images, variation in the task definition, or human raters’ interpretations and downstream goals [becker2019variability, joskowicz2019inter]. Failure to take this ambiguity into account can affect downstream analysis, diagnosis, and treatment.
Recent work tackles these issues separately. In-context learning methods generalize to unseen medical image segmentation tasks, employing an input context or prompt to guide inference [butoi2023universeg, wang2023seggpt, wang2023images]. These methods are deterministic and predict a single segmentation for a given input image and task.
Separately, stochastic or probabilistic segmentation methods output multiple plausible segmentations at inference, reflecting the task uncertainty [baumgartner2019phiseg, kohl2018probabilistic, monteiro2020stochastic]. Each such model is trained for a specific task, and can only output multiple plausible segmentations at inference for that task. Training or fine-tuning a model for a new task requires technical expertise and computational resources that are often unavailable in biomedical settings.
@@ 2 Related Work
Biomedical segmentation is a widely-studied problem, with recent methods dominated by UNet-like architectures [ronneberger_u-net_2015, badrinarayanan2017segnet, isensee2021nnu]. These models tackle a wide variety of tasks, such as different anatomical regions, different structures to segment within a region, different image modalities, and different image settings. With most methods, a new model has to be trained or fine-tuned for each combination of these. Additionally, most models don’t take into account image ambiguity, and provide a single deterministic output.
Multiple Predictions. Uncertainty estimation can help users decide how much faith to put in a segmentation [czolbe2021segmentation] and guide downstream tasks. Uncertainty is often categorized into aleatoric, uncertainty in the data, and epistemic, uncertainty in the model [der2009aleatory, kendall2017uncertainties]. In this work, we focus on aleatoric uncertainty. Medical images are also heteroscedastic in that the degree of uncertainty varies across the image.
Different strategies exist to capture uncertainty. One can assign a probability to each pixel [kendall2015bayesian, hong2021hypernet, islam2021spatially, larrazabal2021maximum], or use contour strategies and difference loss functions to predict the largest and smallest plausible segmentations [lambert2023triadnet, ye2023confidence]. These strategies however do not capture the correlations across pixels. To address this, some methods generate multiple plausible label maps given an image [monteiro2020stochastic, baumgartner2019phiseg, kohl2018probabilistic, kohl2019hierarchical, wolleb_diffusion_2021]. To achieve this, one can directly model pixel correlations, such as through a multivariate Gaussian distribution (with low rank) covariance [monteiro2020stochastic], or more complex distributions [bhat2022generalized]. Alternatively, various frameworks combine potentially hierarchical representations for UNet-like architectures with variational auto-encoders [kohl2018probabilistic, baumgartner2019phiseg, kohl2019hierarchical]. More recently, diffusion models have been used for ensembling [wolleb_diffusion_2021] or to produce stochastic segmentations [rahman2023ambiguous, zbinden2023stochastic]. Some methods explicitly model the different annotators to capture ambiguity [hu2023inter, nichyporuk2022rethinking, schmidt2023probabilistic, tanno2019learning]. But these methods do not apply to our framework where the number of annotators and their characteristics are unknown.
@@
In-context Learning. Few-Shot frameworks use a small set of examples to generalize to new tasks [ding2023few, li2022prototypical, nguyen2019feature, shen2022q, pandey2023robust, seo2022task, zhang2019canet], sometimes by fine-tuning an existing pretrained model [finn2017model, nichol2018first, vinyals2016matching, snell2017prototypical]. In-context learning segmentation methods (ICL) use a small set of examples directly as input to infer label maps for a task [butoi2023universeg, balavzevic2023towards, kim2023universal, wang2023seggpt, wang2023seggpt, kim2023universal]. This enables them to generalize to new tasks. For example, UniverSeg uses an enhanced UNet-based architecture to generalize to medical image segmentation tasks unseen during training [butoi2023universeg]. We build on these ideas to enable segmentation of new tasks without the need to re-train, but expand this paradigm to model stochastic segmentations.
Test Time Augmentation. The test-time augmentation (TTA) strategy uses perturbations of a test input and ensembles the resulting predictions. Existing TTA frameworks model accuracy [shanmugam2021better, melanieTAAL, sinha2019variational, kim2021task], robustness [cohen2023boosting], and estimates of uncertainty [matsunaga2017image, ayhan2022test]. Test-time augmentation has been applied to diverse anatomies and modalities including brain MRI and retinal fundus [amiri2020two, ayhan2022test, huang2021style, moshkov2020test, ResNetTTAColon, wang2019automatic]. Prior work has formalized the variance of a model’s predictions over a set of input transformations as capturing aleatoric uncertainty [TTA2019aleatoric, wang2019automatic, ayhan2022test].
@@ 3.1.1 Neural Network
We use a convolutional architecture focused on interacting representations of sets of flexible sizes using a modified version of the usual UNet structure [ronneberger_u-net_2015].
@@
SetBlock. We introduce a new operation called SetBlock, which interacts the candidate representations $U=\{u_{i}\}_{i=1}^{K}$, with the context representations $V=\{v_{i}\}_{i=1}^{S}$, illustrated in Figure 4. We use the CrossBlock [butoi2023universeg] as a building block for this new layer. The $\text{CrossBlock}(u,V)\rightarrow(u^{\prime},V^{\prime})$ compares an image representation $u$ to a context set representation $V$ through convolutional and averaging operations, and outputs a new image representation $u^{\prime}$ and a new set representation $V^{\prime}$ (Figure 3). $\text{SetBlock}(U,V)\rightarrow(U^{\prime},V^{\prime})$ builds on CrossBlock and performs a set to set interaction of the entries of $U$ and $V$:
@@
where $||$ is the concatenation operation along the feature dimension. The CrossBlock interacts the context representation with the mean candidate. The $\textit{Conv}_{m}$ step communicates this result to all candidate representation. $\textit{Conv}_{u}$ and $\textit{Conv}_{v}$ then update all representations. All convolution operations include a non-linear activation function.
@@ 3.1.2 Best candidate Loss
where $y_{r}^{t}$ is a segmentation from rater $r$, and $\mathcal{L}_{Dice}$ is a weighted sum of soft Dice loss [milletari2016v] and categorical crossentropy. By only back-propagating through the best prediction among $K$ candidates, the network is encouraged to produce diverse solutions [charpiat2008automatic, guzman2012multiple, kirillov2023segment, li2018interactive].
@@ 3.1.3 Training Data
We employ a large dataset of single- and multi-rater segmentations across diverse biomedical domains. We then use data augmentation [butoi2023universeg], as described in B.3.
@@ 4.1 Data
Single-Annotator Data. For single annotator data, we build on MegaMedical used in recent publications [butoi2023universeg, wong2023scribbleprompt] and employ a collection of 73 datasets, of different public biomedical domains and different modalities [ACDC, PanDental, cDemris, CAMUS, CoNSeP, CHAOS2021Challenge, OCTA500, Rose, IDRID, DRIVE, CheXplanation, LUNA, SCD, SpineWeb, VerSe, WBC, BBBC003, butoi2023universeg, ssTEM, MMOTU, ISBI_EM, PanNuke, ISIC, LiTS, NCI-ISBI, KiTS, AMOS, CHAOSdata2019, SegTHOR, BTCV, Promise12, I2CVB, Word, MSD, DukeLiver, CT2US, CT_ORG, SegThy, NerveUS, BRATS, MCIC, ISLES, WMH, BrainDevelopment, OASIS-data, PPMI, LGGFlair, PAXRay, HipXRay, ToothSeg, BUID, BUSIS]. MegaMedical spans a variety of anatomies and modalities, including brain MRI, cardiac ultrasound, thoracic CT and dental X-ray. We also use synthetic data involving simulated shapes, intensities, and image artifacts [butoi2023universeg, hoffmann2021synthmorph]. The single-annotator datasets used for out-of-domain (O.D.) testing are: PanDental [PanDental], WBC [WBC], SCD [SCD], ACDC [ACDC], and SpineWeb [SpineWeb].
Multi-Annotator Data. For multi-annotator I.D. data, we use four datasets from Qubiq [qubiq]: Brain Growth, Brain Lesions, Pancreas Lesions, and Kidney. We also simulate a multi-rater dataset consisting of random shapes (blobs). For the O.D. multi-annotator data we use four datasets. One contains hippocampus segmentation maps on brain MRIs from a large hospital. We crop the volumes around the hippocampus [kohl2018probabilistic] to focus on the areas where the raters disagree. The second is a publicly available lung nodule dataset, LIDC-IDRI [armato2011lung]. This dataset is notable for the substantial inter-rater variability. It contains 1018 thoracic CT scans, each annotated by 4 annotators from a pool of 12 annotators. Finally, we also use retinal fundus images, STARE [STARE], annotated by 2 raters, and prostate data from the MICCAI 2021 QUBIQ challenge [qubiq], annotated by 6 raters on two tasks. Single and multi-annotator combined, our O.D. group contains 20 tasks unseen at training time (some datasets have several tasks).
@@ 4.2 Evaluation
In the multi-annotator setting, we evaluate using both best candidate Dice score, also called maximum Dice score, as well as Generalized Energy Distance (GED)[bellemare2017cramer, salimans2018improving, szekely2013energy]. GED is commonly used in the stochastic segmentation literature to asses the difference between the distribution of predictions and the distribution of annotations [kohl2018probabilistic, baumgartner2019phiseg, monteiro2020stochastic, rahman2023ambiguous, zbinden2023stochastic]. GED has limitations, such as rewarding excessive prediction diversity [rahman2023ambiguous]. Let $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ be the set of annotations, GED is defined as:
@@
[…]
@@ 4.3 Benchmarks
In-Context Methods. We compare to deterministic frameworks that can leverage a context set: a few-shot method, SENet [roy2020SEnet], and two in-context learning (ICL) methods, UniverSeg [butoi2023universeg] and SegGPT [wang2023seggpt]. We train UniverSeg and SENet with the same data split strategies and the same sets of augmentation transforms as for Tyche. For SegGPT, we use the public model, trained on a mix of natural and medical images. Figure 13 in the Supplemental Material shows that UniverSeg trained with additional data outperforms its public version.
Stochastic Upper Bounds. We compare to task-specialized probabilistic segmentation methods that are trained-on and perform well on specific datasets. We independently train Probabilistic UNet [kohl2018probabilistic], PhiSeg[baumgartner2019phiseg] and CIDM, a recent diffusion network [rahman2023ambiguous], on each of the 20 held-out tasks. For each task, we train three model variants: no augmentation, weak augmentation, and as much augmentation as for the Tyche targets. For each benchmark variant, we train on a O.D. development split and select the model that performs best on the corresponding O.D. validation split. We then compare these benchmarks to Tyche on the held-out O.D. test splits.
@@
Interactive Segmentation Methods. We compare to two interactive methods: SAM [kirillov2023segment] and SAM-Med2D [cheng_sam-med2d_2023]. These methods can provide multiple segmentations, but, unlike Tyche, require human interaction, which is outside our scope. SAM also has a functionality to segment all elements in an image, however less optimized for medical imaging. We assume that the SAM-based models have access to the same information as the ICL methods: several image-segmentation pairs as context to guide the segmentation task. We fine-tune SAM using our I.D. development datasets. To replace the human interaction, we provide a bounding box, the average context label map, and 10 clicks, 5 positive and 5 negative as input. With SAM-Med2D, we use a bounding box, and several positive and negative clicks as input. For both SAM and SAM-Med2D, we generate clicks and bounding box from the average context label map.
@@
SegGPT $0.10\pm 0.01$ $0.68\pm 0.01$ $0.94\pm 0.01$ $89\pm 0.03 0.0$$02\pm 0.01$
In-Context SENet $68\pm 0.01 0.0$$00\pm 0.00$ $0.83\pm 0.02$ $0.83\pm 0.02$ $0.30\pm 0.03$
@@ 4.4 Experiments
Inference Setting. We use a fixed context of 16 image-segmentation pairs, because existing in-context learning systems show minimal improvements beyond this size [butoi2023universeg]. Because there is variability in performance depending on the context sampled, we sample 5 different context sets for each datapoint and average performance. Similarly, for the stochastic upper bounds and interactive methods, we do 5 rounds of sampling $K_{i}$ samples.
@@ 5.3 Inference Runtime
UniverSeg $96.62\pm 0.61$ 1.21.2M
SegGPT $2857.19\pm 4.38$ 370370M
SENet $14.91\pm 0.21$ 0.890.89M
FT-SAM $1036.75\pm 4.61$ 9494M
SAM-Med2D $188.8\pm 7.58$ 9191M
PhiSeg $11.35\pm 0.672$ 21.121.1M
ProbaUNet $8.44\pm 0.46$ 55M
CIDM $1.7\times 10^{5}\pm 2748$ 85.685.6M
Tyche-IS $8.57\pm 2.626 1.21.2$M
Tyche-TS $18.09\pm 0.61$ 1.71.7M
@@ B.1 Medical Image Data
Megamedical. We build on the dataset collection proposed in [butoi2023universeg, wong2023scribbleprompt], using the similar preprocessing methods. The complete list of datasets is presented Table 10.
@@
Synthetic Data. We use a set of synthetic tasks to enhance the performance of our networks, similar to [butoi2023universeg]. Some examples are shown in Figure 11.
@@ B.2 Training of Tyche-TS
translate[$\in[0,0.1]$
Random Affine 0.25 scale[$\in[0.9,1.1]$
brightness[$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast[$\in[0.5,1.5]$
@@
translate[$\in[0,0.2]$
Random Affine 0.5 scale[$\in[0.8,1.1]$
brightness[$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast[$\in[0.8,1.2]$